# Training a Scientific Reasoning Model for Chemistry

### v1

@@ 1 Introduction
The dominant approach to improve the accuracy of large language models (LLMs) in recent years has been to scale pre-training corpora size and pre-training compute budget [1, 2, 3, 4]. Partly driven by the finite availability of pre-training data, however, attention has shifted towards alternative scaling dimensions. Such dimensions include strategies such as majority voting [5, 6], “budget-forcing” [7], and test-time training [8], which attempt to scale inference compute. Broadly, reasoning models attempt to improve performance emitting their thought process before arriving at an answer. Early approaches in this vein attempted to elicit reasoning behavior through chain-of-thought (CoT) prompting [9, 10]. More recently, however, reasoning behavior has been demonstrated to emerge through reinforcement learning (RL) post-training, without the need for CoT-style prompting.
@@
RL post-training represents a shift of focus from pre-training data to problems with verifiable rewards. Solutions to such problems can be checked for correctness, allowing the model to generate new, verifiable outputs during learning, explore the space of solutions, and overcome limits imposed by fixed data resources. Multiple works have demonstrated the potential of this approach, particularly in the domains of mathematics and programming. These include both closed-source models [11, 12], and more recently, a large number of open-source models [13, 14, 15, 16, 17, 18, 19].
@@
Scientific domains may be particularly well suited for reasoning models because, as in mathematics and programming, it is often straightforward to assess the quality of a solution, but much more difficult to generate a solution. For example, we may be able to accurately measure the solubility of a given molecule, yet designing a molecule that achieves a desired solubility can be a significant challenge. These “inverse problems” are common in many areas of the physical sciences [20, 21, 22, 23, 24]. More broadly, the scientific method is grounded in structured reasoning: formulating a hypothesis based on observation, evaluating the logical implications of the hypothesis based on experiment, and refining the hypothesis based on analysis of the results of experiment. Science often involves cognitive strategies such as breaking problems into subproblems, responding to failures, or reasoning backwards from desired outcomes, which are strategies also exhibited by reasoning models [25]. However, despite the conceptual alignment between science and reasoning models, there is still relatively little work on scientific reasoning models, aside from benchmarks on multiple choice questions [26, 27, 28].
@@
In this work, we focus on chemistry, with tasks centered on designing, completing, modifying, or synthesizing molecules. This setting is a good demonstration for scientific reasoning models. First, molecules can be represented in text in the SMILES format [29, 30, 31], avoiding the complexities of training a modality-specific encoder. Second, text-based representations of molecules are short relative to modalities in materials science and biology such as nucleotide sequences or CIF files. Third, generating and editing molecules is a critical application, where novel compounds may lead to meaningful clinical and commercial advancements.
@@
We demonstrate the efficacy of reasoning models in chemical tasks by introducing ether0, a novel model that reasons in natural language and outputs molecular structures as SMILES strings. On the chemical reasoning tasks under consideration, ether0 outperforms frontier LLMs, human experts, and models trained for general chemistry. To efficiently train our model, we utilize a series of optimizations over vanilla RL, including distillation of reasoning behaviors, a dynamic curriculum, and initializing RL with distillation from expert models. We further analyze ether0’s data efficiency, failure modes, and reasoning behavior to understand the utility of a reasoning in solving chemistry problems.
@@
Reasoning models are characterized by an attempt to impart system 2-type decision-making [32] to LLMs. Early efforts to this affect include chain-of-thought (CoT) [9], zero-shot CoT [10], and Tree of Thought (ToT) [33] which seek to elicit reasoning by modifying LLM prompts. Later attempts make use of process-level supervision to provide feedback on individual reasoning steps [34, 35, 36]. Most recently, a number of reasoning models have been released [12, 37, 13, 11, 17, 38, 39] using large-scale reinforcement learning via Group Relative Policy Optimization (GRPO)  [40] or inference time scaling [41, 7].
@@
While frontier reasoning models have been evaluated on chemistry tasks [18, 37, 11], the vast majority of these benchmarks have consisted of chemical “knowledge” tasks rather than chemical reasoning tasks [13]. While datasets such as GPQA-D [28], MMLU [26], MMLU-Pro [27], OlympiadArena [42], and Humanity’s last exam [43] assess chemistry knowledge, they do not assess the model’s ability to perform sophisticated chemical reasoning tasks such as retrosynthesis and proposing new structures. While many works have evaluated non-reasoning LLMs on chemical reasoning tasks [44, 45, 46], used LLMs as components for chemical tasks [47, 48, 49], or investigated CoT-style prompting strategies [50, 51], to the best of our knowledge there have been no attempts to directly train reasoning models to perform chemical reasoning tasks using large-scale reinforcement learning. In terms of other scientific domains, OmniScience [52] targets general science applications through distillation on reasoning traces. Med-R1 [53] applies GRPO to medical vision-language tasks, using reinforcement learning to improve generalization and clinically grounded behavior across multi-modal diagnostic reasoning tasks. BioReason [54] integrates a DNA foundation model with an LLM and combines supervised fine-tuning and GRPO to enable interpretable, multi-step genomic reasoning.
@@ 2 Chemical Reasoning Tasks
We construct a dataset of 640,730 chemical reasoning problems, comprising 18 different tasks. The answer to all problems is either a molecule or a reaction. Many tasks are broken down into subtasks. For example, in the solubility editing task, one subtask is to increase solubility without changing the molecular scaffold, and another is to change it without affecting specific functional groups. Table 1 summarizes all problems in our dataset, and Section C.1 provides full details on the dataset provenance as well as the construction of each task. Molecules are represented in the question and expected answer as SMILES, which encodes the molecular graph or chemical reaction as ASCII text [55].
Table 1: Breakdown of verifiable reward training tasks. Machine learning (ML) model verifier means a trained predictive model (i.e. regressing solubility or predicting reaction products). MCQ refers to multiple choice questions. Templates are unique phrasings of the questions in each category. Data source name is a short-name (see citations for complete attribution). Not a sum, because multiple choice property questions use the same template. Also does a “reasonable molecule” check.
@@
Solubility edit 3 115977 ML model[56], code 15 ChEMBL[57]
IUPAC name 1 74994 code 10 COCONUT[58, 59]
SMILES completion 1 74990 code 10 COCONUT[58, 59]
Molecular formula 1 18738 code 10 COCONUT[58, 59]
Functional group 1 74562 code 6 ChEMBL[57]
Elucidation 1 74164 code 10 COCONUT[58, 59]
Retrosynthesis 1 67252 ML model[60], Bloom filter[61] 8 -
Reaction prediction 1 61205 code 10 ORD[62, 63]
Molecule caption 1 54148 code 8 LlaSMol[64]
Safety 11 5687 MCQ 8 Pubchem[65]
Scent 180 4240 MCQ 8 pyFUME[66, 67, 68, 69, 70, 71, 72, 73, 74, 75]
Blood-brain barrier 2 2064 MCQ 8 BBB[76]
Receptor binding 150 1663 MCQ 8 EveBio[77]
ADME 12 1030 MCQ 8 Fang ADME[78]
Aqueous solubility 2 464 MCQ 8 AqSolDB[79]
LD50 2 342 MCQ 8 Pubchem [65]
pKa 4 336 MCQ 8 IUPAC[80]
Photoswitches 1 23 MCQ 8 Photoswitches[81]
@@
We strove to only use synthesized molecules when constructing our dataset, in contrast to previous work in cheminformatics based on “hypothetical” molecules [82]. Thus, all the questions and answers are based on the result of physical experiments. Full reward function implementation details are provided in Section C.2. In addition to the criteria listed, tasks marked with also check that the proposed molecules are plausibly synthesizable by fragmentation into rings and local groups (details in Section C.2).
Solubility edit: Modify a given molecule to increase or decrease aqueous solubility ($\log S$). Subtasks impose additional constraints enforcing similarity to the input molecule. The $\log S$ objective is computed using KDESol [56] and constraints are evaluated using RDKit [83] and exmol [84].
@@
Molecular formula: Propose a molecule given a molecular formula in Hill notation [85]. Verified with RDKit.
@@
Elucidation: Determine the chemical structure of a molecule found in an organism given its molecular formula and background information on the organism. Since the problem is underdetermined, we consider any answer to be correct if the proposed molecule has a Tanimoto similarity (ECFP4 [86]) of at least 0.7 to the ground truth. Verified with RDKit.
Retrosynthesis: Provide a single-step reaction to produce the given target molecule. The reactants must all be purchasable molecules (determined by manufacturer catalogs in a Bloom filter [61]), and the product of the proposed reaction must match the target molecule predicted using the Molecular Transformer model [60].
Reaction prediction: Given a chemical reaction, predict the major product. Verify exact molecule match with RDKit. Molecular caption: Given a textual description of a molecule, produce the SMILES of the molecule. This task uses data from Yu et. al [64], which itself comes from PubChem [87, 88, 89]. Verified with RDKit.
Multiple choice questions: Predict or modify properties of a molecule, for which no accurate oracle exists. Instead, multiple options are presented, and the model is expected to select the one that has been experimentally determined to satisfy the criterion. Verified by string matching.
@@ 3 Background
As in prior work [90, 13], we use SFT to initialize a policy for RL (Equation S1). If the demonstration dataset $\mathcal{D}_{\mathrm{demo}}$ is itself from another policy $\pi^{\prime}$, this can also be considered a form of expert iteration [91, 92] or knowledge distillation [93].
@@
While SFT can be used to warm-start the policy, we rely heavily on online reinforcement learning to improve our models. In particular, we use Group Relative Policy Optimization (GRPO) [40].
@@
where $\pi_{\theta}$ is the policy being optimized, $\pi_{\theta_{\mathrm{old}}}$ is the policy from which we sampled rollouts, and $\pi_{\mathrm{ref}}$ is a reference policy. $\mathrm{clip}$ is the standard PPO clip function [94]:
@@ 4 Training
In this section, we describe a method to train a large language model to reason about and answer the problems detailed in Section 2. We utilize a multi-stage training procedure, consisting of alternating phases of (a) distillation [93] and (b) GRPO [40, 13]. At a high level, the stages are: (1) Supervised fine-tuning on long chain-of-thought reasoning sequences; (2) Task-specific “specialist” GRPO; (3) Distillation of specialist models into an all-task “generalist” model; and (4) Generalist GRPO. Using a family of task-specific reasoning models to generate synthetic data for a generalist model has been recently demonstrated to be an effective strategy in other domains [95, 96].
Unless otherwise stated, our policies are trained from Mistral-Small-24B-Instruct-2501 [97].
To simplify formatting of the model output, we introduce four new tokens to the base model’s vocabulary to demarcate reasoning and answering boundaries. During distillation and RL, these tokens are used to respectively format and validate sequences with the following structure:
@@ 4.1 Long CoT Supervised Fine-Tuning
The SFT sequences are first generated by prompting DeepSeek-R1 with a subset of the training dataset, with a maximum token budget of 8192 tokens. To remove low-quality responses, we enforce the following criteria: (1) each sequence ends with an answer enclosed in XML tags; (2) the answer is valid SMILES/SMIRKS; and (3) passes an LLM-based check for relevant reasoning (Section B.1). We considered rejecting responses with incorrect answers, but R1’s success rate is below 1% for many tasks. Our goal during SFT is to find a good pre-RL initialization, not necessarily to maximize accuracy, and prior work [13, 25] suggests that SFT even on inaccurate reasoning sequences can be sufficient. Therefore, we do not discard sequences that end in incorrect answers.
@@ 4.2 Specialist RL
where format_reward is 1 if the format is met and 0 otherwise; accuracy_reward is 1 if the answer satisfies the problem (Section 2) and 0 otherwise. The only exception is the specialist trained on molecular formula, functional group, and elucidation, which uses a softer accuracy_reward: if the desired molecular formula is met but other constraints are not, then 0.5 is returned.
Note that RL allows to bootstrap new behaviors not present in the SFT traces. An example of this can be shown in  Section F.1.
4.3 Advantage-Based Curriculum
The GRPO advantage reduces to zero on groups in which all elements achieve the same reward. Besides the KL term, these “trivial” groups do not contribute to the policy gradient, and their fraction of the batch $f_{T}$ can reach $90\%$ during training. DAPO [98] tackles this by discarding trivial groups and resampling problems, requiring $\sim(1-f_{T})^{-1}\times$ as many sampling attempts per batch.
We instead use a heuristic: if a problem results in a non-trivial group from the current policy, it is added to a curriculum buffer. At each training iteration, a fraction ($\epsilon_{\mathrm{cur}}$) of the batch is selected from the buffer instead of the dataset. Since these problems were recently non-trivial, we expect a lower $f_{T}$ than the rest of the dataset. If a buffer problem becomes trivial, it is removed from the curriculum. This method can reduce $f_{T}$ with no additional computational cost, demonstrated in Section F.3. A similar method has been previously employed in the offline setting, using reward variance [99].

### v2

@@ 1 Introduction
The dominant approach to improve the accuracy of large language models (LLMs) in recent years has been to scale pre-training corpora size and pre-training compute budget radford2019language; brown2020language; kaplan2020scaling; 2022_Hoffmann. Partly driven by the finite availability of pre-training data, however, attention has shifted towards alternative scaling dimensions. Such dimensions include strategies such as majority voting 2024_Brown; 2024_Aviary, “budget-forcing” muennighoff2025s1, and test-time training akyurek2024surprising, which attempt to scale inference compute. Broadly, reasoning models attempt to improve performance emitting their thought process before arriving at an answer. Early approaches in this vein attempted to elicit reasoning behavior through chain-of-thought (CoT) prompting 2022_Wei; 2022_Kojima. More recently, however, reasoning behavior has been demonstrated to emerge through reinforcement learning (RL) post-training, without the need for CoT-style prompting.
@@
RL post-training represents a shift of focus from pre-training data to problems with verifiable rewards. Solutions to such problems can be checked for correctness, allowing the model to generate new, verifiable outputs during learning, explore the space of solutions, and overcome limits imposed by fixed data resources. Multiple works have demonstrated the potential of this approach, particularly in the domains of mathematics and programming. These include both closed-source models anthropic2025claude; jaech2024openai, and more recently, a large number of open-source models 2025_Guo; 2025_Kimi; deepscaler2025; deepcoder2025; 2025_Bercovich; 2025_Akter; 2025_Blakeman.
@@
Scientific domains may be particularly well suited for reasoning models because, as in mathematics and programming, it is often straightforward to assess the quality of a solution, but much more difficult to generate a solution. For example, we may be able to measure the solubility of a given molecule, yet designing a molecule with a desired solubility can be a significant challenge. These “inverse problems” are common in many areas of the physical sciences arridge2019solving; ongie2020deep; siahkoohi2023reliable; zheng2025inversebench; 2021_Lesnic. More broadly, the scientific method is grounded in structured reasoning: formulating a hypothesis based on observation, evaluating the logical implications of the hypothesis based on experiment, and refining the hypothesis based on analysis of the results of experiment. Science often involves cognitive strategies such as breaking problems into subproblems, responding to failures, or reasoning backwards from desired outcomes, which are strategies also exhibited by reasoning models gandhi2025cognitive. However, despite the conceptual alignment between science and reasoning models, there is still relatively little work on scientific reasoning models, aside from benchmarks on multiple choice questions 2021_Hendrycks; 2024_Wang; 2024_Rein.
@@
In this work, we focus on chemistry, with tasks centered on designing, completing, modifying, or synthesizing molecules. This setting is a good demonstration for scientific reasoning models. First, molecules can be represented in text in the SMILES format 1988_Weininger; 1989_Weininger; 1990_Weininger, avoiding the complexities of training a modality-specific encoder. Second, text-based representations of molecules are short relative to modalities in materials science and biology such as nucleotide sequences or CIF files. Third, generating and editing molecules is a critical application, where novel compounds may lead to meaningful clinical and commercial advancements.
We demonstrate the efficacy of reasoning models in chemical tasks by introducing ether0, a novel model that reasons in natural language and outputs molecular structures as SMILES strings. On the chemical reasoning tasks under consideration, ether0 outperforms frontier LLMs, human experts, and models trained for general chemistry. Moreover, ether0 supports key stages of the drug discovery pipeline: it can generate candidates during hit discovery, it enables molecule editing in hit-to-lead to improve potency, selectivity, or physicochemical properties, and it contributes to lead optimization by refining compounds to enhance efficacy, reduce toxicity, and improve ADMET profiles, all while being aware of synthesizability.
@@
To efficiently train our model, we utilize a series of optimizations over vanilla RL, including distillation of reasoning behaviors, a dynamic curriculum, and initializing RL with distillation from expert models. We further analyze ether0’s data efficiency, failure modes, and reasoning behavior to understand the utility of a reasoning in solving chemistry problems.
@@
Reasoning models are characterized by an attempt to impart system 2-type decision-making 2011_Kahneman to LLMs. Early efforts to this affect include chain-of-thought (CoT) 2022_Wei, zero-shot CoT 2022_Kojima, and Tree of Thought (ToT) 2023_Yao which seek to elicit reasoning by modifying LLM prompts. Later attempts make use of process-level supervision to provide feedback on individual reasoning steps 2023_Hao; 2024_Gao; 2024_Lightman. Most recently, a number of reasoning models have been released jaech2024openai; o3; 2025_Guo; anthropic2025claude; 2025_Bercovich; 2024_Qin; 2024_Huang using large-scale reinforcement learning via Group Relative Policy Optimization (GRPO)  2024_Shao or inference time scaling 2025_Huang; muennighoff2025s1.
@@
While frontier reasoning models have been evaluated on chemistry tasks 2025_Akter; o3; anthropic2025claude, the vast majority of these benchmarks have consisted of chemical “knowledge” tasks rather than chemical reasoning tasks 2025_Guo. While datasets such as GPQA-D 2024_Rein, MMLU 2021_Hendrycks, MMLU-Pro 2024_Wang, OlympiadArena 2024_Arena, and Humanity’s last exam 2025_Phan assess chemistry knowledge, they do not assess the model’s ability to perform sophisticated chemical reasoning tasks such as retrosynthesis and proposing new structures. While many works have evaluated non-reasoning LLMs on chemical reasoning tasks 2023_Guo; 2024_Li; 2024_Feng, used LLMs as components for chemical tasks 2023_Boiko; 2024_Bran; 2025_Bran, or investigated CoT-style prompting strategies 2024_Ouyang; 2024_Jang, to the best of our knowledge there have been no attempts to directly train reasoning models to perform chemical reasoning tasks using large-scale reinforcement learning. In terms of other scientific domains, OmniScience prabhakar2025omniscience targets general science applications through distillation on reasoning traces. Med-R1 2025_Lai applies GRPO to medical vision-language tasks, using reinforcement learning to improve generalization and clinically grounded behavior across multi-modal diagnostic reasoning tasks. BioReason 2025_Fallahpour integrates a DNA foundation model with an LLM and combines supervised fine-tuning and GRPO to enable interpretable, multi-step genomic reasoning.
@@ 2 Chemical Reasoning Tasks
We construct a dataset of 640,730 chemical reasoning problems, comprising 18 different tasks. Molecules are represented in the question and expected answer as SMILES, which encodes the molecular graph or chemical reaction as ASCII text 2018_Kochev. The answers are all either a molecule or a reaction. Many tasks are broken down into subtasks. For example, in the solubility editing task, one subtask is to increase solubility without changing the molecular scaffold, and another is to change it without affecting specific functional groups. Table 1 summarizes all problems in our dataset, and Section C.2 provides full details on the dataset provenance as well as the construction of each task.
Table 1: Breakdown of verifiable reward training tasks. ML model verifier: trained predictive model; MCQ: multiple-choice questions. Templates: unique phrasings per category. Data source: short name (see citations for full attribution). Not a sum; multiple-choice property questions share templates. Also performs a “reasonable molecule” check.
@@
Solubility edit 3 115977 ML model2024_Ramos, code 15 ChEMBLzdrazil2024chembl
IUPAC name 1 74994 code 10 COCONUTsorokina2021coconut; chandrasekhar2025coconut
SMILES completion 1 74990 code 10 COCONUTsorokina2021coconut; chandrasekhar2025coconut
Molecular formula 1 18738 code 10 COCONUTsorokina2021coconut; chandrasekhar2025coconut
Functional group 1 74562 code 6 ChEMBLzdrazil2024chembl
Elucidation 1 74164 code 10 COCONUTsorokina2021coconut; chandrasekhar2025coconut
Retrosynthesis 1 67252 ML model2019_Schwaller, Bloom filtermedina2023bloom 8 -
Reaction prediction 1 61205 code 10 ORD2021_Kearnes; 2023_Mercado
Molecule caption 1 54148 code 8 LlaSMol2024_Yu
Safety 11 5687 MCQ 8 Pubchemkim2023pubchem
Scent 180 4240 MCQ 8 pyFUMEKeller_2012; Bushdid_2014; Garg_2017; Keller_2016; WYSOCKI_1989; Ravia_2020; Sharma_2021; Snitz_2013; Snitz_2019; hamel2024pyrfume
Blood-brain barrier 2 2064 MCQ 8 BBBmeng2021curated
Receptor binding 150 1663 MCQ 8 EveBioEvE_Bio_Pharmome_R3_2025
ADME 12 1030 MCQ 8 Fang ADMEfang2023prospective
Aqueous solubility 2 464 MCQ 8 AqSolDB2019_Sorkun
LD50 2 342 MCQ 8 Pubchem kim2023pubchem
pKa 4 336 MCQ 8 IUPACZheng_IUPAC_pKa_2025
Photoswitches 1 23 MCQ 8 Photoswitches2022_Griffiths
@@
We strove to only use synthesized molecules when constructing our dataset, in contrast to previous work in cheminformatics based on “hypothetical” molecules ruddigkeit2012enumeration. Thus, all the questions and answers are based on the result of physical experiments. Full reward function implementation details are provided in Section C.3. In addition to the criteria listed, tasks marked with also check that the proposed molecules are plausibly synthesizable by fragmentation into rings and local groups (details in Section C.3).
Solubility edit: Modify a given molecule to increase or decrease aqueous solubility ($\log S$). Subtasks impose additional constraints enforcing similarity to the input molecule. The $\log S$ objective is computed using KDESol 2024_Ramos and constraints are evaluated using RDKit rdkit and exmol 2022_Wellawatte.
@@
Molecular formula: Propose a molecule given a molecular formula in Hill notation 1900_Hill. Verified with RDKit.
@@
Elucidation: Determine the chemical structure of a molecule found in an organism given its molecular formula and background information on the organism. Since the problem is underdetermined, we consider any answer to be correct if the proposed molecule has a Tanimoto similarity (ECFP4 rogers2010extended) of at least 0.7 to the ground truth. Verified with RDKit.
Retrosynthesis: Provide a single-step reaction to produce the given target molecule. The reactants must all be purchasable molecules (determined by manufacturer catalogs in a Bloom filter medina2023bloom), and the product of the proposed reaction must match the target molecule predicted using the Molecular Transformer model 2019_Schwaller.
Reaction prediction: Given a chemical reaction, predict the major product. Verify exact molecule match with RDKit.
Molecular caption: Given a textual description of a molecule, produce the SMILES of the molecule. This task uses data from Yu et. al 2024_Yu, which itself comes from PubChem 2016_Kim; 2019_Kim; 2023_Kim. Verified with RDKit.
Multiple choice questions: Predict or modify properties of a molecule, for which no accurate oracle exists. Instead, multiple options are presented, and the model is expected to select the one that has been experimentally determined to satisfy the criterion. Verified by string matching. See Section C.1.
@@ 3 Background
As in prior work 2022_Ouyang; 2025_Guo, we use SFT to initialize a policy for RL (Equation S1). If the demonstration dataset $\mathcal{D}_{\mathrm{demo}}$ is itself from another policy $\pi^{\prime}$, this can also be considered a form of expert iteration 2017_Anthony; 2024_Havrilla or knowledge distillation 2015_Hinton.
@@
While SFT can be used to warm-start the policy, we rely heavily on online reinforcement learning to improve our models. In particular, we use Group Relative Policy Optimization (GRPO) 2024_Shao.
@@
where $\pi_{\theta}$ is the policy being optimized, $\pi_{\theta_{\mathrm{old}}}$ is the policy from which we sampled rollouts, and $\pi_{\mathrm{ref}}$ is a reference policy. $\mathrm{clip}$ is the standard PPO clip function 2017_Schulman:
@@ 4 Training
In this section, we describe a method to train a large language model to reason about and answer the problems detailed in Section 2. We utilize a multi-stage training procedure, consisting of alternating phases of (a) distillation 2015_Hinton and (b) GRPO 2024_Shao; 2025_Guo. At a high level, the stages are: (1) Supervised fine-tuning on long chain-of-thought reasoning sequences; (2) Task-specific “specialist” GRPO; (3) Distillation of specialist models into an all-task “generalist” model; and (4) Generalist GRPO. Using a family of task-specific reasoning models to generate synthetic data for a generalist model has been recently demonstrated to be an effective strategy in other domains 2024_Abdin; 2025_DeepSeekV3.
Unless otherwise stated, our policies are trained from Mistral-Small-24B-Instruct-2501 2025_Mistral. To simplify formatting of the model output, we introduce four new tokens to the base model’s vocabulary to demarcate reasoning and answering boundaries. During distillation and RL, these tokens are used to respectively format and validate sequences with the following structure:
@@ 4.1 Long CoT Supervised Fine-Tuning
The SFT sequences are first generated by prompting DeepSeek-R1 with a subset of the training dataset, with a maximum token budget of 8192 tokens. To remove low-quality responses, we enforce the following criteria: (1) each sequence ends with an answer enclosed in XML tags; (2) the answer is valid SMILES/SMIRKS; and (3) passes an LLM-based check for relevant reasoning (Section B.1). We considered rejecting responses with incorrect answers, but R1’s success rate is below 1% for many tasks. Our goal during SFT is to find a good pre-RL initialization, not necessarily to maximize accuracy, and prior work 2025_Guo; gandhi2025cognitive suggests that SFT even on inaccurate reasoning sequences can be sufficient. Therefore, we do not discard sequences that end in incorrect answers.
@@ 4.2 Specialist RL
where format_reward is 1 if the format is met and 0 otherwise; accuracy_reward is 1 if the answer satisfies the problem (Section 2) and 0 otherwise. The only exception is the specialist trained on molecular formula, functional group, and elucidation, which uses a softer accuracy_reward: if the desired molecular formula is met but other constraints are not, then 0.5 is returned. Note that RL allows to bootstrap new behaviors not present in the SFT traces. An example of this can be shown in  Section F.1.
4.2.1 Advantage-Based Curriculum
The GRPO advantage reduces to zero on groups in which all elements achieve the same reward. Besides the KL term, these “trivial” groups do not contribute to the policy gradient, and their fraction of the batch $f_{T}$ can reach $90\%$ during training. DAPO 2025_Yao tackles this by discarding trivial groups and resampling problems, requiring $\sim(1-f_{T})^{-1}\times$ as many sampling attempts per batch.
We instead use a heuristic: if a problem results in a non-trivial group from the current policy, it is added to a curriculum buffer. At each training iteration, a fraction ($\epsilon_{\mathrm{cur}}$) of the batch is selected from the buffer instead of the dataset. Since these problems were recently non-trivial, we expect a lower $f_{T}$ than the rest of the dataset. If a buffer problem becomes trivial, it is removed from the curriculum. This method can reduce $f_{T}$ with no additional computational cost, demonstrated in Section F.3. A similar method has been previously employed in the offline setting, using reward variance 2025_WangSingleExample.